O_{2}(c^{1}Σ_{u}^{−}–b^{1}Σ_{g}^{+}) BAND EMISSION IN THE TERRESTRIAL NIGHTGLOW

\underline{DAVID L. HUESTIS}, PHILIP C. COSBY, and TOM G. SLANGER, Molecular Physics Laboratory, SRI International, Menlo Park, CA 94025.

Understanding of the spectroscopy of molecular oxygen has benefited from many investigations of the absorption and emission spectra from planetary atmospheres, a number of which we have described previously at this symposium. In the process of line-by-line assignment of “sky spectra” recorded by astronomers using the Keck I and II telescopes, we have discovered a new O_{2} band system in the 380-450 nm region of the Earth’s nightglow spectrum.

Sharp isolated lines are assigned to the single Q-branch of O_{2}(c^{1}Σ_{u}^{−},v′=5-11) radiating to O_{2}(b^{1}Σ_{g}^{+},v′′=0-2), with intensities that follow the J′(J′+1)(2J′+1)exp[- B′J′(J′+1)/kT] formula for rotational coupling with ^{1}Π states suggested by Wilkinson and Mulliken [1] for the corresponding a′-X transition in N_{2}. Assignment of lines up to J′=32 allows for improvement of centrifugal distortion constants for the O_{2}(c) state [2,3], while the measured positions of the low J lines agree with those calculated from known term energies [2-4].

The individual lines in the 9-1 c-b band are as intense as any lines in the Herzberg I (A-X) or Chamberlain (A′-a) bands at wavelengths longer than 365 nm. Surprisingly, Herzberg II (c-X) band emissions are much weaker than expected from previous studies of the terrestrial [5,6] and venusian nightglows [7,8].

[1] P. G. Wilkinson and R. S. Mulliken, J. Chem. Phys. 31, 674 (1959).

[2] D. A. Ramsay, Can. J. Phys. 64, 171 (1986).

[3] K. Yoshino, J. R. Esmond, W. H. Parkinson, A. P. Thorne, R.C.M. Learner, and G. Cox, J. Chem. Phys. 111, 2960 (1999).

[4] H. D. Babcock and L. Herzberg, Astrophys. J. 108, 167 (1948).

[5] T. G. Slanger and D. L. Huestis, J. Geophys. Res. 86, 3551 (1981); 88, 4137 (1983).

[6] J. Stegman and D. P. Murtach, Planet. Space Sci. 36, 927 (1988).

[7] V. A. Krasnopolsky, A. A. Krysko, V. N. Rogachev, and V. A. Parshev, Cosmic Res. 13, 687 (1977).

[8] T. G. Slanger, P. C. Cosby, D. L. Huestis, and T. A. Bida, Science 291, 463 (2001).